OpenStudy (anonymous):

tan x/1 + sec x + 1 + sec x/ tan x = 2 csc x

OpenStudy (bahrom7893):

can u post the parenthesis in proper places?

OpenStudy (anonymous):

(tan x/1) + sec x + 1 + (sec x/ tan x) = 2 csc x

OpenStudy (bahrom7893):

tan x/1 is just tangent.. i thik u got the parenthesis wrong. And it would help if u stopped reposting the same thing

OpenStudy (anonymous):

tan x = sin x/ cos x what do you mean its just tangent???

OpenStudy (bahrom7893):

Tan = Sin/Cos Sec = 1/Cos Csc=1/Sin

OpenStudy (anonymous):

i know so how do i prove that that equals 2 csc x

OpenStudy (bahrom7893):

(tan x/1) + sec x + 1 + (sec x/ tan x) = 2 csc x Are u sure u meant (tan x/1) because that would be just tan x. Didn't u mean: tanx/(1+sec x) or something?

OpenStudy (anonymous):

someone answered earlier ago but they're not responding

OpenStudy (bahrom7893):

where?

OpenStudy (anonymous):

and he gave a really good and thorough answer

OpenStudy (anonymous):

satellite73

OpenStudy (bahrom7893):

That doesn't help, i know satellite, where's the question. If he already answered, then why do u want another answer?

OpenStudy (anonymous):

b/c i didnt understand this one step..

OpenStudy (bahrom7893):

link me to the question

OpenStudy (anonymous):

there.. i wrote down what i didnt get

OpenStudy (anonymous):

still confusion?